If 1.63 x 10^-4 g of helium dissolves in 100.0 g of waer, what is the concentration in parts per million?
1
Expert's answer
2013-01-16T11:38:01-0500
C = N of He /N of H2O

N of H2O = m/Mw * 6.02*10^23 = 100.0/18 *6.02*10^23 =33.44 *10^23

N of He =m/Mw * 6.02*10^23 = (1.63 x 10^-4 / 4) *6.02*10^23= 0.000245 *10^23

If you have 0.000245 *10^23 atoms of He
per 33.44 *10^23 molecules of H2O
You will have X atoms of He per 1000000

X = 7.34 parts per million